OpenStudy (anonymous):

I will give out medals. Rachael has 6 cups of milk for her cereal. She puts 3/4 cup in her cereal bowl each day. How long will Rachel's milk last? Express your answer in simplest form.

OpenStudy (anonymous):

Im not sure if im correct, but im guessing you just count how many times 3/4 goes into 6

OpenStudy (anonymous):

I think Rachel will be able to use her milk 4 times . but im not sure !

OpenStudy (anonymous):

What we need to do is figure out have many times 3/4 goes into 6. We do this by dividing 6 by 3/4: \[\frac{6}{\frac{3}{4}}\] Instead of doing this division we can make it easier by knowing that dividing by a fraction is the same as multiplying by its reciprocal. If we do this we get: \[\frac{6}{1} \times \frac{4}{3} = \frac{24}{3} = 8\]

OpenStudy (anonymous):

If you want a quick check you can convert 3/4 to 0.75 and divide: \[\frac{6}{0.75}=8\]

OpenStudy (anonymous):

thx